OpenStudy (anonymous):

A business evaluates a proposed venture as follows. It stands to make a profit of $10,000 with probability 3/20, to make a profit of $5000 with probability 9/20, to break even with probability 5/20, and to lose $5000 with probability 3/20. What is the expected profit in dollars?

OpenStudy (kropot72):

Let the profit (loss) be P, and the expected profit be E(P): \[\large E(P)=(10,000\times\frac{3}{20})+(5000\times\frac{9}{20})+(0\times\frac{5}{20})-(5000\times\frac{3}{20})\]